The motion detector used in a physics lab sends and receives 30 kHz ultrasonic pulses. A pulse goes out, reflects off the object being measured, and returns to the detector. The lab temperature is 20 ∘C.

Part A) What is the wavelength of the waves emitted by the motion detector? Express your answer in millimeters.

Part B) How long does it take for a pulse that reflects off an object 3.0 m away to make a round trip? Express your answer in milliseconds.
Answer #1

(a) The speed of sound in air at 20 degrees c is 343m/s.

Wavelength = c/f

…Where…

c = speed of sound in m per second (343m/s)

f = frequency in hertz (30,000)

…therefore…

wavelength = 343/30000 = 0.0114 m

(b)

Round trip to object 3m away = 2*3 = 6m

Time (s) = Distance (m)/c

= 6/343

=0.0175 seconds